# Andreev bound states in Josephson junctions of semi-Dirac semimetals

### 31 Dec 2023 v1

@@ Abstract
We consider a Josephson junction built with the two-dimensional semi-Dirac semimetal, which features a hybrid of linear and quadratic dispersion around a nodal point. We model the weak link between the two superconducting regions by a Dirac delta potential because it mimics the thin-barrier limit of a superconductor-barrier-superconductor configuration. Assuming a homogeneous pairing in each region, we set up the BdG formalism for electronlike and holelike quasiparticles propagating along the quadratic-in-momentum dispersion direction. This allows us to compute the discrete bound-state energy spectrum $\varepsilon$ of the subgap Andreev states localized at the junction. In contrast with the Josephson effect investigated for propagation along linearly dispersing directions, we find a pair of doubly degenerate Andreev bound states. Using the dependence of $\varepsilon$ on the superconducting phase difference $\phi$, we compute the variation of Josephson current as a function of $\phi$.
I Introduction
In a configuration consisting of two superconductors coupled together by a weak link between them, a dissipationless current flows across the junction in equilibrium [1, 2, 3], which is dubbed as the Josephson current $I_{J}$. $I_{J}$ is a single-valued and $2\pi$-periodic function the phase difference $\phi$ of the pair potential in the two superconductors. In such a set-up, the Andreev surface states of the two superconductors hybridize to form Andreev bound states (ABSs) at the junction. These states are the dominant contributors to the Josephson current through the junction [4, 5, 6], because the contributions from the excited states in the continuum are negligible. There is an extensive literature devoted to the study of such Josephson effects in two-dimensional (2d) and three-dimensional (3d) materials like graphene and Weyl-like semimetals [5, 7, 8, 9, 10, 11, 12, 13], where the weak link is a tunneling barrier. In other words, the two superconducting regions (abbreviated by “S”) are weakly coupled by a middle region consisting of the normal (i.e., non-superconducting) phase (abbreviated by “N”) of a semimetal. Two alternate configurations include S-N-S [5, 7] and S-B-S (where “B” indicates a potential barrier in the N region) [8, 9, 11, 12, 13] junctions. While the superconductivity is induced via proximity-effect by placing a conventional s-wave superconductor on top of the corresponding electrode [14], the tunnel barrier can be created by applying a gate voltage $V_{0}$ across N. A schematic representation of the S-B-S set-up is illustrated in Fig. 1(a).
@@ I Introduction
In this paper, we consider the S-B-S configuration in the thin-barrier limit, constructed with the semi-Dirac semimetal, by approximating the barrier by a Dirac delta function [6, 30, 31]. In our set-up, the weak link is represented by the thin-barrier limit of an S-B-S junction, which is defined by $L\ll\xi$, where $L$ is the barrier thickness and $\xi$ is the superconducting coherence length. We employ the scattering matrix approach for the associated Bogoliubov–de Gennes (BdG) Hamiltonian, which has been one of the standard techniques used extensively to determine the conductance of an N-S junction [32, 33, 34, 35, 36], with the appropriate generalization applicable for the S-B-S junctions.
The propagation direction of the quasipaticles/quasiholes is taken to be parallel/antiparallel to the $x$-axis. We denote the transverse dimension of the junction by $W$, where $W$ is assumed to be large enough to impose periodic boundary conditions along the $z$-directiom. In the short-barrier regime, the main contribution to the Josephson current comes from the subgap Andreev states [4, 5, 6], because the contributions from the excited states in the continuum are suppressed by a factor of $L/\xi$. We compute the energies of the ABSs in the thin-barrier-limit and determine the resulting the Josephson current.
The paper is organized as follows. In Sec. II, we describe the low-energy effective Hamiltonian of the semi-Dirac semimetal in the normal phase, and show its eigenvalues and eigenfunctions. In Secc III, the BdG Hamiltonian necessary to describe the S-B-S junction is shown, along with the expressions for electronlike and holelike wavefunctions. This is followed by Sec. IV, where the ABS energy values are derived and the Josephson current is computed numerically. Finally, we end with a summary and outlook in Sec. V.
II Semi-Dirac semimetal
@@
$\mathcal{H}(\mathbf{k})=\frac{\hbar^{2}\,k_{x}^{2}}{2\,m}\,\sigma_{x}+\hbar\,v\,k_{z}\,\sigma_{z},$ (1)
where $m$ is the effective mass parameter along the $x$-axis, $v$ is the Fermi velocity along the $z$-axis, and $\sigma_{x}$ and $\sigma_{z}$ are two of the three Pauli matrices. In order to simplify the notations, we define the dimensionless momenta
@@
The tight-binding Hamiltonian [15], from which the above low-energy continuum Hamiltonian has been obtained, consists of a honeycomb lattice comprising two sublattice sites labelled as A and B (analogous to the case for graphene). Hence, there exists a pair of valleys (at the $\mathbf{K}$ and $\mathbf{K}^{\prime}$ corners of the Brillouin zone) with opposite chiralities, with $\mathcal{H}(\mathbf{k})$ representing the states in the vicinity of the valley with positive chirality (i.e., $\mathbf{K}$ point of the Brillouin zone).
The energy eigenvalues are then given by
$\mathcal{E}=s\,\varepsilon_{0}\,\sqrt{K_{x}^{4}+K_{y}^{2}}\,,\quad\varepsilon_{0}=\frac{p^{2}}{2\,m}=2\,p\,v\,,\quad s=\pm\,,$ (3)
@@
as shown in Fig. 1(b). Here the “$+$” and “--” signs, as usual, refer to the conduction band (i.e., upper band with positive energy eigenvalue) and valence band (i.e., lower band with negative energy eigenvalue), respectively. Henceforth, we set $\varepsilon_{0}$ to unity for uncluttering of notations, which just implies that all our energies are scaled in units of $\varepsilon_{0}$. Furthermore, we will take $\hbar=1$ except occasionally for clarity.
The set of two orthonormal eigenvectors is given by
$\Psi^{s}(\mathbf{k})=\frac{1}{\sqrt{K_{x}^{4}+\left(s\,E+K_{z}\right)^{2}}}\left[s\,E+K_{z}\quad K_{x}^{2}\right]^{T},\quad E=\sqrt{K_{x}^{4}+K_{z}^{2}}\,.$ (4)
@@
In addition to the propagating plane wave solutions, there are also evanescent waves [16, 38, 39, 40] present because of the fact that, for a given value of the transverse (to the propagation direction) momentum $K_{z}$ with $|K_{z}|\leq|E|$, the relation $K_{x}^{4}=E^{2}-K_{z}^{2}$ leads to the four solutions $K_{x}=\pm\ \mathcal{K}_{x}$ and $K_{x}=\pm\,\tilde{\mathcal{K}}_{x}$, where $\mathcal{K}_{x}=\left(E^{2}-K_{z}^{2}\right)^{1/4}$ and $\tilde{\mathcal{K}}_{x}=i\,\mathcal{K}_{x}$. If the Fermi energy cuts the bands at energy $E$, then for propagation along the $x$-direction, the corresponding “right-moving” plane waves will have the factor $e^{i\,\text{sgn}(E)\,\tilde{\mathcal{K}}_{x}\,x}$ — this just implies that if the propagating quasiparticles are occupying the upper(lower) band, then they have a positive(negative) group velocity.
@@
We note that the “extra” solutions involving the evanescent waves (for example, the ones with momentum ±$\pm\,\tilde{\mathcal{K}}_{x}$) do not exist when the propagation direction is taken to be along the $z$-axis, as the dispersion is linear-in-momentum along that direction. As a result, we expect a richer structure of the ABSs when we analyze the problem of junctions encountered for transmission along the $x$-direction, which features a nonlinear dependence on the momentum.
III Josephson junction
@@
In order to represent the superconducting phases of the Josephson junction [cf. Fig. 1(a)], we need to define the superconducting pair potential in each region. The time-reversal operator interchanges the valleys in a graphene-like bandstructure with pseudospin-1/2, where each valley is represented by a Hamiltonian constructed out of the Pauli matrices. Because of the valley degeneracy, it suffices to consider one of the two possible sets, thus leading to a $4\times 4$ BdG Hamiltonian. Adopting a homogeneous approximation, we follow the construction in Refs. [32, 8] such that the pair potential can modelled as
@@
representing BCS-like Cooper pairing in the spin-singlet s-wave channel. The phases of the superconducting order parameter in the two regions are given by $\phi_{1}$ and $\phi_{1}+\phi$, such that the phase difference is $\phi$. Since the final expressions for the ABSs and the Josephson current depend only on $\phi$, we set $\phi_{1}=0$ without any loss of generality. Due to the presence of the Delta function potential barrier between the two superconductors, we need to consider the potential energy function
@@
A few important points need to be remembered here in order to contrast our scenario with the linear-in-momentum dispersion cases. Although the thin-barrier limit is equivalent to a Dirac delta potential, we do not have any constraint on the derivatives of the wavefunction across the junctions when the dispersion is linear, implying that the standard delta function potential approximation for thin barriers cannot be taken from the outset [8]. For those situations, we need to start with a finite normal state region (rather than a Dirac delta function), obtain the equations from the boundary conditions at the S-B and B-S junctions, and finally impose the appropriate limits while computing the final solutions [5, 8, 10, 11, 13]. However, for a quadratic dispersion, we can use the Dirac delta approximation from the start because here we have a constraint on the first order derivatives (with respect to the position coordinate along the propagation direction) of the wavefunction across the junction, analogous to the tunneling problem involving a Schrödinger particle.
The BdG Hamiltonian can be constructed as
$H=\sum_{\mathbf{r}}\Psi^{\dagger}(\mathbf{r})\,H_{\text{BdG}}(\mathbf{r})\,\Psi(\mathbf{r}),\quad\Psi(\mathbf{r})=\begin{bmatrix}c_{A+}(\mathbf{r})&c_{B+}(\mathbf{r})&c_{A-}^{\dagger}(\mathbf{r})&-c_{B-}^{\dagger}(\mathbf{r})\end{bmatrix}^{T},$ (8)
$H_{\text{BdG}}(\mathbf{r})=\begin{bmatrix}\mathcal{H}(\mathbf{K}\rightarrow\bm{-}i\,\nabla_{\mathbf{r}})-E_{F}+V(x)&\Delta(x)\\ \Delta^{\dagger}(x)&E_{F}-V(x)-\mathcal{H}(\mathbf{K}\rightarrow\bm{-}i\,\nabla_{\mathbf{r}})\\ \end{bmatrix},$ (11)
where where $\mathbf{r}=(x,\,y)$ is the position vector and the indices $\pm$ on the fermionic operators label the two valleys of opposite chiralities.
Here we demarcate the left superconducting region as “L” and the right superconducting region as “R”, with the Delta function barrier being the weak link in the middle. The electron-like and hole-like BdG quasiparticles are obtained from the eigenvalue equation
@@
If $\psi_{N}(\mathbf{K})$ denotes an eigenfunction of $\mathcal{H}(\mathbf{K})$, then the electronlike and holelike eigenfunctions of $H_{\text{BdG}}(\mathbf{r})$ are given by the expressions [41]
$\psi_{e}(\mathbf{K})=\begin{bmatrix}\psi_{N}(\mathbf{K})&\frac{\left(\varepsilon-\Omega\right)\,e^{-i\,\varphi}}{\Delta_{0}}\,\psi_{N}(\mathbf{K})\end{bmatrix}\text{ and }\psi_{h}(\mathbf{K})=\begin{bmatrix}\psi_{N}(\mathbf{K})&\frac{\left(\varepsilon+\Omega\right)\,e^{-i\,\varphi}}{\Delta_{0}}\,\psi_{N}(\mathbf{K})\end{bmatrix},$ (15)
respectively, where
@@
    $\psi_{L}(\mathbf{r},K_{z})$ $=a_{l}\,\psi_{e}(-\kappa_{x}^{e},K_{z},0)\,e^{-i\,\kappa_{x}^{e}\,x+K_{z}\,z}+b_{l}\,\psi_{h}(-\kappa_{x}^{h},K_{z},0)\,e^{-i\,\kappa_{x}^{h}\,x+K_{z}\,z}$
    $\quad+c_{l}\,\psi_{e}(-\tilde{\kappa}_{x}^{e},K_{z},0)\,e^{-i\,\tilde{\kappa}_{x}^{e}\,x+K_{z}\,z}+d_{l}\,\psi_{h}(\tilde{\kappa}_{x}^{h},K_{z},0)\,e^{i\,\tilde{\kappa}_{x}^{h}\,x+K_{z}\,z}\,,$ (19)
@@
    $\psi_{e}(K_{x},K_{z},\varphi)$ $\simeq\left[e^{i\,\beta}\left(E_{F}+K_{z}\right)\quad e^{i\,\beta}\,K_{x}^{2}\quad e^{-i\,\varphi}\left(E_{F}+K_{z}\right)\quad e^{-i\,\varphi}\,K_{x}^{2}\right]^{T},$
    $\psi_{h}(K_{x},K_{z},\varphi)$ $\simeq\left[K_{z}-E_{F}\quad K_{x}^{2}\quad e^{i\,\beta-i\,\varphi}\left(K_{z}-E_{F}\right)\quad e^{i\,\beta-i\,\varphi}\,K_{x}^{2}\right]^{T},$ (20)
@@
    $\psi_{L}(\mathbf{r},K_{z})$ $=a_{r}\,\psi_{e}(\kappa_{x}^{e},K_{z},\phi)\,e^{i\,\kappa_{x}^{e}\,x+K_{z}\,z}+b_{r}\,\psi_{h}(\kappa_{x}^{h},K_{z},\phi)\,e^{i\,\kappa_{x}^{h}\,x+K_{z}\,z}$
    $\quad+c_{r}\,\psi_{e}(\tilde{\kappa}_{x}^{e},K_{z},\phi)\,e^{i\,\tilde{\kappa}_{x}^{e}\,x+K_{z}\,z}+d_{r}\,\psi_{h}(-\tilde{\kappa}_{x}^{h},K_{z},\phi)\,e^{-i\,\tilde{\kappa}_{x}^{h}\,x+K_{z}\,z}\,.$ (22)
@@
Imposing the continuity of the wavefunction and the constraint on its first order derivatives (with respect to $x$) across the junction located at $x=0$, we get the following equations:
$\psi_{L}(0,z,K_{z})=\psi_{R}(0,z,K_{z})\text{ and }\partial_{x}\psi_{L}(x,z,K_{z})\big{|}_{x=0}-\partial_{x}\psi_{R}(x,z,K_{z})\big{|}_{x=0}=V_{0}\,\psi_{L}(0,z,K_{z})\,.$ (23)

### 08 Apr 2024 v2

@@ Abstract
We consider a Josephson junction built with the two-dimensional semi-Dirac semimetal, which features a hybrid of linear and quadratic dispersion around a nodal point. We model the weak link between the two superconducting regions by a Dirac delta potential because it mimics the thin-barrier-limit of a superconductor-barrier-superconductor configuration. Assuming a homogeneous pairing in each region, we set up the BdG formalism for electronlike and holelike quasiparticles propagating along the quadratic-in-momentum dispersion direction. This allows us to compute the discrete bound-state energy spectrum $\varepsilon$ of the subgap Andreev states localized at the junction. In contrast with the Josephson effect investigated for propagation along linearly dispersing directions, we find a pair of doubly degenerate Andreev bound states. Using the dependence of $\varepsilon$ on the superconducting phase difference $\phi$, we compute the variation of Josephson current as a function of $\phi$.
1 Introduction
In a configuration consisting of two superconductors coupled together by a weak link between them, a dissipationless current flows across the junction in equilibrium [1, 2, 3], which is dubbed as the Josephson current $I_{J}$. $I_{J}$ is a single-valued and $2\pi$-periodic function in the phase difference $\phi$ of the pair potential of the two superconductors. In such a set-up, the Andreev surface states of the two superconductors hybridize to form Andreev bound states (ABSs) at the junction. These states are the dominant contributors to the Josephson current through the junction [4, 5, 6, 7], because the contributions from the excited states in the continuum are negligible. There is an extensive literature devoted to the study of such Josephson effects in two-dimensional (2d) and three-dimensional (3d) materials like graphene and Weyl-like semimetals [5, 8, 9, 10, 11, 12, 13, 14], where the weak link is a tunneling barrier. In other words, the two superconducting regions (each abbreviated by “S”) are weakly coupled by a middle region consisting of the normal (i.e., non-superconducting) phase (abbreviated by “N”) of a semimetal. Two alternate configurations include the S-N-S [5, 8] and the S-B-S (where “B” indicates a potential barrier in the N region) [9, 10, 12, 13, 14] junctions. While the superconductivity is induced via proximity-effect by placing a conventional s-wave superconductor on top of the corresponding electrode [15], the tunnel barrier can be created by applying a gate voltage $V_{0}$ across N. A schematic representation of the S-B-S set-up is illustrated in Fig. 1(a).
@@ 1 Introduction
In this paper, we consider the S-B-S configuration in the thin-barrier-limit, constructed with the semi-Dirac semimetal, by approximating the barrier by a Dirac delta function [6, 32, 33]. In our set-up, the weak link is represented by the thin-barrier-limit of an S-B-S junction, which is defined by $L\ll\xi$, where $L$ is the barrier thickness and $\xi$ is the superconducting coherence length. We employ the scattering matrix approach for the associated Bogoliubov–de Gennes (BdG) Hamiltonian, which has been one of the standard techniques used extensively to determine the conductance of an N-S junction [34, 35, 36, 37, 38], with the appropriate generalization applicable for the S-B-S junctions.
The propagation direction of the quasiparticles/quasiholes is taken to be parallel/antiparallel to the $x$-axis. We denote the transverse dimension of the junction by $W$, where $W$ is assumed to be large enough to impose periodic boundary conditions along the $z$-directiom. In the short-barrier regime, the main contribution to the Josephson current comes from the subgap Andreev states [4, 5, 6], because the contributions from the excited states in the continuum are suppressed by a factor of $L/\xi$. We compute the energies of the ABSs in the thin-barrier-limit, and determine the resulting Josephson current.
The paper is organized as follows. In Sec. 2, we describe the low-energy effective Hamiltonian of the semi-Dirac semimetal in the normal phase, and show its eigenvalues and eigenfunctions. In Sec 3, the BdG Hamiltonian, necessary to describe the S-B-S junction, is shown along with the expressions for the electronlike and holelike wavefunctions. This is followed by Sec. 4, where the ABS energy values are derived and the Josephson current is computed numerically. Finally, we end with a summary and outlook in Sec. 5.
2 Semi-Dirac semimetal
@@
$\mathcal{H}(\mathbf{k})=\frac{\hbar^{2}\,k_{x}^{2}}{2\,m}\,\sigma_{x}+\hbar\,v\,k_{z}\,\sigma_{z}\,.$ (1)
Here, $m$ is the effective mass parameter along the $x$-axis, $v$ is the Fermi velocity along the $z$-axis, and $\sigma_{x}$ and $\sigma_{z}$ are two of the three Pauli matrices. In order to simplify the notations, we define the dimensionless momenta
@@
The tight-binding Hamiltonian [16], from which the above low-energy continuum Hamiltonian has been obtained, consists of a honeycomb lattice comprising two sublattice sites labelled as A and B (analogous to the case of graphene). Hence, there exists a pair of valleys at the two inequivalent K points/corners of the Brillouin zone, which we denote by ${}_{+}$ and ${}_{-}$. The Hamiltonian $\mathcal{H}(\mathbf{k})$ here represents the states in the vicinity of the valley located at ${}_{+}$.
The energy eigenvalues of $\mathcal{H}(\mathbf{k})$ are given by
$\mathcal{E}=s\,\varepsilon_{0}\,\sqrt{K_{x}^{4}+K_{z}^{2}}\,,\quad\varepsilon_{0}=\frac{p^{2}}{2\,m}=2\,p\,v\,,\quad s=\pm\,,$ (3)
@@
as shown in Fig. 1(b). Here, the “$+$” and “--” signs, as usual, refer to the conduction band (i.e., the upper band with the positive energy eigenvalue) and the valence band (i.e., the lower band with the negative energy eigenvalue), respectively. Henceforth, we set $\varepsilon_{0}$ to unity for uncluttering of notations, which just implies that all our energies are scaled in units of $\varepsilon_{0}$. Furthermore, we will set $\hbar=1$, except occasionally, when we retain it for the sake of clarity.
A set of two orthonormal eigenvectors is given by
$\Psi^{s}(\mathbf{k})=\frac{1}{\sqrt{K_{x}^{4}+\left(s\,E+K_{z}\right)^{2}}}\left[s\,E+K_{z}\qquad K_{x}^{2}\right]^{T},$
$E=\sqrt{K_{x}^{4}+K_{z}^{2}}\,.$ (4)
@@
For a given value of the transverse (to the propagation direction) momentum $K_{z}$ with $|K_{z}|\leq|E|$, the relation $K_{x}^{4}=E^{2}-K_{z}^{2}$ leads to the four solutions $K_{x}=\pm\ \mathcal{K}_{x}$ and $K_{x}=\pm\,\tilde{\mathcal{K}}_{x}$, where $\mathcal{K}_{x}=\left(E^{2}-K_{z}^{2}\right)^{1/4}$ and $\tilde{\mathcal{K}}_{x}=i\,\mathcal{K}_{x}$. Consequently, in addition to the propagating plane wave solutions, there are also evanescent waves present [17, 40, 41, 42]. If the Fermi energy cuts the bands at energy $E$, then for propagation along the $x$-direction, the corresponding “right-moving” plane waves will have the factor $e^{i\,\text{sgn}(E)\,\tilde{\mathcal{K}}_{x}\,x}$ — this just implies that if the propagating quasiparticles are occupying the upper(lower) band, then they have a positive(negative) group velocity.
@@
We note that the “extra” solutions involving the evanescent waves (i.e., the ones with the momenta ±$\pm\,\tilde{\mathcal{K}}_{x}$) do not exist when the propagation direction is taken to be along the $z$-axis, as the dispersion is linear-in-momentum along that direction. As a result, we expect a richer structure of the ABSs when we analyze the problem of junctions encountered for transmission along the $x$-direction, which features a nonlinear dependence on the momentum.
3 Josephson junction
@@
In order to represent the superconducting phases across the Josephson junction [cf. Fig. 1(a)], we need to define the superconducting pair potential in each region. The time-reversal operator interchanges the valleys in a graphene-like bandstructure (which has pseudospin-1/2 quasiparticles), where each valley is represented by a Hamiltonian constructed out of the Pauli matrices. Because of the valley degeneracy, it suffices to consider one of the two possible sets, thus leading to a $4\times 4$ BdG Hamiltonian. Adopting a homogeneous approximation, we follow the construction in Refs. [34, 9], such that the pair potential can be modelled as
@@
representing BCS-like Cooper pairing in the spin-singlet s-wave channel. The phases of the superconducting order parameter in the two regions are given by $\phi_{1}$ and $\phi_{1}+\phi$, respectively, such that the phase difference is $\phi$. Since the final expressions for the ABSs and the Josephson current depend only on $\phi$, we set $\phi_{1}=0$, without any loss of generality. Due to the presence of the Delta function potential barrier between the two superconductors, we need to consider the potential energy function
@@
In order to contrast our scenario with the linear-in-momentum dispersion cases, a few important points need to be remembered here. Although the thin-barrier-limit is equivalent to a Dirac delta potential, we do not have any constraint on the derivatives of the wavefunction across the junctions when the dispersion is linear, implying that the standard delta function potential approximation for thin barriers cannot be taken from the outset [9]. For those situations, we need to start with a finite normal state region (rather than a Dirac delta function), obtain the equations from the boundary conditions at the S-B and B-S junctions, and finally impose the appropriate limits while computing the final solutions [5, 9, 11, 12, 14]. However, for a quadratic dispersion, we can use the Dirac delta approximation from the start, because here we have a constraint on the first order derivatives (with respect to the position coordinate along the propagation direction) of the wavefunction across the junction, analogous to the tunneling problem involving a Schrödinger particle.
The BdG Hamiltonian can be constructed as [34, 9]
$H=\sum_{\mathbf{r}}\Psi^{\dagger}(\mathbf{r})\,H_{\text{BdG}}(\mathbf{r})\,\Psi(\mathbf{r}),\quad\Psi(\mathbf{r})=\begin{bmatrix}c_{A+}(\mathbf{r})&&c_{B+}(\mathbf{r})&&c_{A-}^{\dagger}(\mathbf{r})&&-c_{B-}^{\dagger}(\mathbf{r})\end{bmatrix}^{T},$ (8)
$H_{\text{BdG}}(\mathbf{r})=\begin{bmatrix}\mathcal{H}(\mathbf{K}\rightarrow\bm{-}i\,\nabla_{\mathbf{r}})-E_{F}+V(x)&&\Delta(x)\\ \\ \Delta^{\dagger}(x)&&E_{F}-V(x)-\mathcal{H}(\mathbf{K}\rightarrow\bm{-}i\,\nabla_{\mathbf{r}})\\ \end{bmatrix},$ (12)
where $\mathbf{r}=(x,\,z)$ is the position vector, and the indices $\pm$ on the fermionic operators label the two valleys.
Here we demarcate the left superconducting region as “L” and the right superconducting region as “R”, with the delta function barrier being the weak link in the middle. The electronlike and holelike BdG quasiparticles are obtained from the eigenvalue equation
@@
If $\psi_{N}(\mathbf{K})$ denotes an eigenfunction of $\mathcal{H}(\mathbf{K})$, then the electronlike and holelike eigenfunctions of $H_{\text{BdG}}(\mathbf{r})$ are of the forms $\psi_{e}(\mathbf{K},\varphi)\,e^{i\,\mathbf{K}\cdot\mathbf{r}}$ and $\psi_{h}(\mathbf{K},\varphi)\,e^{i\,\mathbf{K}\cdot\mathbf{r}}$, respectively, where [43]
$\psi_{e}(\mathbf{K},\varphi)=\begin{bmatrix}\psi_{N}(\mathbf{K})&&&&\frac{\left(\varepsilon-\Omega\right)\,e^{-i\,\varphi}}{\Delta_{0}}\,\psi_{N}(\mathbf{K})\end{bmatrix}\text{ and }$ (15)
$\psi_{h}(\mathbf{K},\varphi)=\begin{bmatrix}\psi_{N}(\mathbf{K})&&&&\frac{\left(\varepsilon+\Omega\right)\,e^{-i\,\varphi}}{\Delta_{0}}\,\psi_{N}(\mathbf{K})\end{bmatrix}.$ (17)
Here,
@@
    $\psi_{L}(\mathbf{r},K_{z})$ $=a_{l}\,\psi_{e}(-\kappa_{x}^{e},K_{z},0)\,e^{-i\,\kappa_{x}^{e}\,x+K_{z}\,z}$
    $\quad+b_{l}\,\psi_{h}(-\kappa_{x}^{h},K_{z},0)\,e^{-i\,\kappa_{x}^{h}\,x+K_{z}\,z}$
    $\quad+c_{l}\,\psi_{e}(-\tilde{\kappa}_{x}^{e},K_{z},0)\,e^{-i\,\tilde{\kappa}_{x}^{e}\,x+K_{z}\,z}$
    $\quad+d_{l}\,\psi_{h}(\tilde{\kappa}_{x}^{h},K_{z},0)\,e^{i\,\tilde{\kappa}_{x}^{h}\,x+K_{z}\,z}\,,$ (21)
@@
    $\psi_{e}(K_{x},K_{z},\varphi)$ $\simeq\left[e^{i\,\beta}\left(E_{F}+K_{z}\right)\qquad e^{i\,\beta}\,K_{x}^{2}\qquad\qquad e^{-i\,\varphi}\left(E_{F}+K_{z}\right)\qquad e^{-i\,\varphi}\,K_{x}^{2}\right]^{T},$
    $\psi_{h}(K_{x},K_{z},\varphi)$ $\simeq\left[K_{z}-E_{F}\qquad\qquad\quad K_{x}^{2}\qquad\qquad e^{i\,\beta-i\,\varphi}\left(K_{z}-E_{F}\right)\qquad e^{i\,\beta-i\,\varphi}\,K_{x}^{2}\right]^{T},$ (22)
@@
    $\psi_{R}(\mathbf{r},K_{z})$ $=a_{r}\,\psi_{e}(\kappa_{x}^{e},K_{z},\phi)\,e^{i\,\kappa_{x}^{e}\,x+K_{z}\,z}$
    $\quad+b_{r}\,\psi_{h}(\kappa_{x}^{h},K_{z},\phi)\,e^{i\,\kappa_{x}^{h}\,x+K_{z}\,z}$
    $\quad+c_{r}\,\psi_{e}(\tilde{\kappa}_{x}^{e},K_{z},\phi)\,e^{i\,\tilde{\kappa}_{x}^{e}\,x+K_{z}\,z}$
    $\quad+d_{r}\,\psi_{h}(-\tilde{\kappa}_{x}^{h},K_{z},\phi)\,e^{-i\,\tilde{\kappa}_{x}^{h}\,x+K_{z}\,z}\,.$ (24)
@@
Imposing the continuity of the wavefunction and the constraint on its first order derivatives (with respect to $x$) across the junction, located at $x=0$, we get the following equations:
$\psi_{L}(0,z,K_{z})=\psi_{R}(0,z,K_{z})\text{ and }$
$\partial_{x}\psi_{R}(x,z,K_{z})\big{|}_{x=0}-\partial_{x}\psi_{L}(x,z,K_{z})\big{|}_{x=0}=V_{0}\,\psi_{L}(0,z,K_{z})\,.$ (25)